1.   A 400.0-N sign hangs from the end of a uniform strut. The strut is 4.0 m long and weighs 600.0 N. The strut is supported by a hinge at the wall and by a cable whose other end is tied to the wall at a point 3.0 m above the left end of the strut. Find the tension in the supporting cable and the force of the hinge on the strut.



1
Expert's answer
2022-01-02T18:24:03-0500

Solution:


The tension in the supporting cable


"\sum\tau_B=0\to600\cdot2+400\cdot4-4T\cdot(3/5)=0\to T=1167\ (N)" . Answer


The force of the hinge on the strut.


"F_h=T\cdot(4/5)=1167\cdot(4/5)=933\ (N)"


"\sum\tau_A =0\to600\cdot2-F_v\cdot4=0\to F_v=300\ (N)"


"F=\sqrt{933^+300^2}=980\ (N)". Answer